OpenStudy (anonymous):

is y=0.3x a exponential decay function

OpenStudy (ganpat):

try it yourself.. draw a graph.. calculate few values of y for given value of x... For instance.. y = 0.3x 1. x=10 y is 3 2. x=20 y is 6 and few more values... Then draw a graph..

OpenStudy (ganpat):

?? Can you continue from where i have stopped ?

OpenStudy (solomonzelman):

\(\large\color{black}{ y=ab^x }\) is an exponential function: We know that: \(\large\color{black}{ b^0 }\) just like any number raise to the zero (asides from \(\large\color{black}{ 0^0 }\) ) is equal to \(\large\color{black}{1 }\). normally when x=0, then b^x is 1, and therefore the y-intercept is "a" your case is just: \(\large\color{black}{ y=b^x }\) (without "a") you know that when x is 1 you multiply times 0.3 once (saying it is just 0.3) and when x is 2 you multiply 0.3 twice, saying it is \(\large\color{black}{ 0.3 \times 0.3 }\) which is equivalent to 0.09. and so on.